# Non-Linear Scoring Model for Translation Quality Evaluation

### v2

@@ 4 Psychophysical and Cognitive Foundations
Empirical findings from these client interviews show that perceived translation quality is not a simple linear function of error count or text length. Instead, human judgment follows deep-seated perceptual and cognitive regularities documented since the nineteenth century [weber1834, fechner1860, boring1950]. This paper draws on two complementary frameworks—classical psychophysics and modern Cognitive Load Theory—to motivate a logarithmic scoring model.
@@ 4.1 Weber–Fechner Law and Logarithmic Perception
Hence, long before today’s quality assessment challenges, experimental psychology had already established a robust theoretical framework: Human judgment converts objective intensities into logarithm-like internal scales. Treating translation errors as discrete ’stimuli’ fits squarely within this tradition and motivates the logarithmic tolerance curve adopted in this paper.
@@ 4.2 Cognitive Load and Accumulated Impact of Errors
In translation quality evaluation, the same principle applies: each identified error acts as a disruption that requires attentional and interpretive resources. One or two errors may be brushed aside, but a buildup of errors, however minor individually, begins to saturate the evaluator’s cognitive capacity. This leads to increased frustration, reduced comprehension, and a degraded overall impression of quality.
Importantly, this degradation does not occur linearly. Just as in perceptual scaling, there is a threshold effect: the cumulative weight of repeated small errors eventually causes a sharp drop in processing fluency. Studies in instructional design and reading comprehension (e.g. [mcw2020]) have documented how disfluencies compound non-linearly and trigger avoidance behaviors or reduced engagement.
Cognitive Load Theory also provides a plausible explanation for the empirical finding that users tend to expect stricter quality in longer documents, even when the overall error density remains stable [waddington1999, waddington2001]. This reflects a protective mechanism: anticipating that accumulated disruptions will erode comprehension, readers subconsciously raise the standard for what they consider acceptable in large-scale content.
Together, these cognitive effects mirror and reinforce the non-linear sensitivity modeled by the Weber–Fechner framework. Although psychophysics focuses on the perceptual scaling of individual stimuli, Cognitive Load Theory models the impact of cumulative disruption. Both lead to the same conclusion: the relationship between perceived translation quality and the raw error count (that is, the objective number of penalty points) is fundamentally non-linear.
@@ 4.3 Perception and Cognition Converge on Logarithmic Scoring
Together, classical psychophysical scaling [fechner1860, stevens1957] and cognitive load mechanics [sweller1988, miller1956] converge in the same functional form: tolerance for errors grows rapidly at first, then slows logarithmically. A calibrated logarithmic curve therefore offers the most parsimonious way to reconcile (i) diminishing marginal impact of individual errors with (ii) the accelerating cognitive cost of cumulative errors. This synthesis underpins the non-linear scoring model proposed in the next section.
@@ 5 Proposed Non-Linear Scoring Model
This paper adopts the terminology introduced in [GladkoffEtAl2024], distinguishing between three classes of scoring models.
  • Raw Score: $100-\alpha\,\mathrm{APT}/\mathrm{EWC}$, with $\alpha$ a unit factor (e.g., $\alpha=1000$ for “per 1,000 words”). This uses the sample’s penalty rate and is not a calibrated/anchored score.
  • Calibrated Linear: This scoring model maps a client’s acceptance range (pass/fail tolerance) onto a common 0–100 quality scale, making results easy to read and compare across jobs.
  • Calibrated Non-Linear: A model that replaces the straight-line rule with a logarithmic tolerance curve. Relative to a linear rule anchored at the reference size, the model is more lenient for short samples and stricter for long ones; total allowance still grows with length, but at a slowing rate.
@@
Therefore, we recommend the non-linear, logarithmic model for all cases where evaluation sample sizes are expected to vary by more than $20\%$ from the reference wordcount (i.e., outside the $\pm 20\%$ fidelity interval defined in §5.4). Both the traditional raw model and the calibrated linear model remain viable tools, but their use should be restricted to scenarios where evaluation sample sizes consistently hover very close to the specific word count for which the models were calibrated. 444‘Calibration’ for the raw score simply means choosing the client’s pass/fail cut-point on the raw penalty-rate scale. Let $r_{\text{thr}}$ be the maximum allowed error rate (e.g., 5 points per 1,000 words). A sample with penalty total $\mathrm{APT}$ and word count $\mathrm{EWC}$ passes if and only if $\mathrm{APT}/\mathrm{EWC}\leq r_{\text{thr}}$, equivalently $\mathrm{APT}\leq r_{\text{thr}}\cdot\mathrm{EWC}$. No calibration mapping to a 0–100 quality scale is involved. Note that some teams report a 0–100 “raw score” by first scaling the penalty rate to a per-1,000-word unit and then computing $100-1000\cdot\mathrm{APT}/\mathrm{EWC}$; however, this rescaling still inherits the limitations of raw scoring. To obtain comparable scores across evaluation sample sizes, a calibrated model (linear near the reference size or non-linear for wider ranges) is required.
This section introduces the mathematical form of the proposed logarithmic model, explains how it can be calibrated, and demonstrates its practical advantages.
By introducing a Non-Linear Calibrated scoring model, this paper extends the MQM scoring apparatus itself, enabling it to represent perceptual tolerance that linear models cannot capture. Because the same logarithmic function adapts seamlessly across evaluation sample sizes, language pairs, and content domains, an MQM-based Non-Linear Calibrated metric becomes congruent with holistic (document-level) expert judgment—turning MQM into a practically universal translation quality evaluation system. [^5] Other scoring models may be devised for specific applications. Importantly, however, an exponential ladder of severity weights is not the same thing as a non-linear scoring model in the sense used here. Severity weights (even exponential ones) change how individual errors are combined within a fixed sample; they do not by themselves make the pass/fail tolerance as a function of sample size non-linear. For example, the ATA certification rubric is an analytic, points-off scheme tailored to short, fixed-length exam translations. Its severity scale is steep (many exponential levels), so higher-severity errors count far more than minor ones, but aggregation is still essentially additive and the exam length is fixed; there is no tolerance curve $E(x)$ that varies with $x$. Similarly, SAE J2450 (automotive) assigns fixed category weights, producing a non-uniform penalty distribution but still summing points linearly within a sample [saeJ2450]. (One could reproduce similar effects by choosing MQM type weights accordingly.) Thus ATA and SAE J2450 introduce severity-level non-uniformities, not the sample-size non-linearity addressed in this paper. They are effective for their narrow contexts, but they do not provide the general, non-linear tolerance function $E(x)$ needed to keep scores comparable when evaluation size varies widely. A related line of work is the LOGIPEM/HOPE metric we developed for comparing MT engines in cases where human reviewers perceive clear quality differences that standard automatic metrics fail to capture [gladkoff-han-2022-hope]. LOGIPEM/HOPE targets human-aligned system ranking at the document level by shaping how evidence from errors and discourse phenomena is aggregated; it is therefore complementary to the present contribution. In particular, LOGIPEM/HOPE addresses what is measured for engine comparison, whereas the non-linear tolerance function $E(x)$ proposed here addresses how pass/fail tolerance should vary with evaluation size. The former improves agreement with expert judgments across systems; the latter keeps scores consistent and comparable as sample length varies.
5.1 Limitations of Linear Extrapolation
@@
Linear scoring models typically assume that acceptable error counts increase proportionally with the sample size. For example, if a client accepts 5 minor errors in a 1000-word sample, the model allows 10 errors in 2000 words, 15 in 3000, etc. However, as shown in Sections 3 and 4, this assumption may be misaligned with the way content users perceive quality.
@@
As noted above, many institutional clients describe their scoring guidelines in terms that sound linear—roughly “about one serious error per page” as a tolerable upper bound. In practice, however, evaluators do not extend that allowance proportionally. When the sample expands to seven pages, they do not simply treat seven serious errors as acceptable; instead, reviewers report that they would typically fail such a sample already at three or four serious errors. In other words, tolerance tightens as the sample grows: the acceptable error density falls with text length. What matters is not a linear accumulation of defects, but how those defects shape the perceived quality of the translation as a whole. This asymmetry shows why linear extrapolation is misleading: it inflates the acceptable error count in longer samples and diminishes the perceived severity of errors in shorter ones.
5.2 Mathematical Model
@@
Throughout, $\ln$ denotes the natural logarithm.
If $x$ is measured in pages with $W$ words/page, replace $b$ by WbW\,b; $a$ is unchanged. This unit invariance ensures figures shown in pages and formulas written in words are consistent. The curve satisfies $E(0)=0$, rises steeply for short texts, and then flattens. This curve shape effectively combines both theoretical expectations and observed data.
To clarify on the form of this function, the intercept-free form $E(x)=a\,\ln(1+bx)$ is deliberately used, which enforces $E(0)=0$ and allows curvature control via $b$, not the Excel spreadsheet “logarithmic trendline” $c+k\ln x$. The latter introduces a free intercept and does not pass through the origin, so it cannot be anchored to calibration points in the same way.
@@ 5.3 Anchoring and Calibration Procedures
  • Anchoring defines the maximum acceptable error limit for a standard amount of text. This tolerance is expressed in penalty points based on a client’s real-world standard, such as, “no more than 5 penalty points per 1,000 words.” (one minor error is typically equivalent to one penalty point).
@@
Feasibility (two‑point). Let $(x_{0},E_{0})$ and $(x_{1},E_{1})$ be two tolerance points with $x_{0}\neq x_{1}$ and $E_{0},E_{1}>0$. Define
@@
  • A second, empirical data point: This point is derived from observing the client’s real-world judgment (e.g., through a survey), such as their intuitive tolerance for a different sample size—for instance, accepting only 2 errors ($E_{1}=2$) in a 250-word text ($x_{1}=250$).
@@
See  Appendix A. Calibration from Two Points (Numerical) for a one-dimensional root-finding procedure in $b$ (e.g., bisection), followed by $a=E_{0}/\ln(1+b\,x_{0})$.
@@ 5.5 Visual comparison across three sample-size ranges
Figure 5 visualizes this critical weakness of the linear model: Once the slope is fixed using a single calibration point, the linear model severely diverges from the actual human judgment at both lower and higher evaluation sample sizes. In contrast, the logarithmic model aligns more closely with observed data regardless of sample size (evaluation wordcount).
The calibrated logarithmic curve is more permissive for short samples and stricter for long ones relative to an anchored linear rule; it increases with length but flattens, avoiding the linear model’s over‑leniency at large sizes.
This not only matches human content users’ expectations, but also reduces misclassification of borderline quality outputs.
@@
Building on the Multi-Range Framework proposed in [GladkoffEtAl2024], this paper distinguishes three regimes of evaluation sample sizes based on their statistical and perceptual characteristics. In the micro-range (fewer than 250 words), evaluation outcomes are highly sensitive to single errors and statistical confidence is very low; thus, risk-based Statistical Quality Control (SQC) is most appropriate.
@@
Ultimately, however, for any operational scenario involving a wide variety of evaluated text lengths, the non-linear logarithmic model is the only option that ensures consistent and reliable accuracy. This is because a single linear extrapolation becomes unreliable as it moves further from its anchor point, whereas a logarithmic model remains accurate across both the meso- and macro-ranges.
@@
Figure 6 summarizes these applicability ranges, illustrating how tolerance curves change shape across evaluation sample size regimes.
@@ 6 Practical Implications
To the extent that error tolerance is not linear — as suggested by both empirical findings and practitioner reports — the adoption of a logarithmic scoring model can reshape translation quality evaluation practices, quality management processes, and software tooling across the industry. Replacing the linearity assumption allows quality assessment to become more accurate, consistent, and aligned with human perception. Looking ahead, such a model might also serve as a reference framework for emerging AI-based document-level evaluations. Current AI systems often provide holistic but qualitative judgments with uncertain completeness; a calibrated non-linear scoring model could, in principle, offer a quantitative reference point for future systems that integrate or adjust their evaluations against established calibrated 0–100 quality scale (client-specific PT/MSV mapping).
@@ 7 Building your non-linear scorecard
Notation. The designations and symbols in this section follow the variable naming used in the ASTM working draft WK46396 (“MQM 2.0: Analytic Translation Quality Evaluation”), which describes the MQM scoring models and their mechanics in detail.
@@ 7.3 Scorecard mapping
This keeps the MQM table (error types, error weights, severities and related multipliers $\rightarrow$ APT) unchanged; only the tolerance is dynamic via (2) and the calibrated mapping stays the same as the linear model’s PT/DPI (Passing Threshold/Defined Passing Interval) presentation.
Figure 7: Example of a real-world non-linear calibrated scoring model for documentation content type (solid yellow curve) plotted against the linear model with 10 allowed penalty points per 1000 words (dashed line). The logarithmic curve is calibrated from two anchor points (not shown in this plot), and therefore represents the actual client tolerance function; the linear line is drawn for comparison only.

### v3

@@ 4 Psychophysical and Cognitive Foundations
Empirical findings from these anonymized institution interviews show that perceived translation quality is not a simple linear function of error count or text length. Instead, human judgment follows deep-seated perceptual and cognitive regularities documented since the nineteenth century [weber1834, fechner1860, boring1950]. This paper draws on two complementary frameworks—classical psychophysics and modern Cognitive Load Theory—to motivate a logarithmic scoring model.
@@ 4.1 Weber–Fechner Law and Logarithmic Perception
Hence, long before today’s quality assessment challenges, experimental psychology had already established a robust theoretical framework: Human judgment converts objective intensities into logarithm-like internal scales. Treating translation errors as discrete ‘stimuli’ fits squarely within this tradition and motivates the logarithmic tolerance curve adopted in this paper.
@@ 4.2 Cognitive Load and Accumulated Impact of Errors
In translation quality evaluation, the same principle applies: each identified error acts as a disruption that requires attentional and interpretive resources. One or two errors may be brushed aside, but a buildup of errors, however minor individually, begins to saturate the evaluator’s cognitive capacity. This leads to increased frustration, reduced comprehension, and a degraded overall impression of quality. Importantly, this degradation does not occur linearly. Just as in perceptual scaling, there is a threshold effect: the cumulative weight of repeated small errors eventually causes a sharp drop in processing fluency. Studies in instructional design and reading comprehension (e.g. [mcw2020]) have documented how disfluencies compound non-linearly and trigger avoidance behaviors or reduced engagement.
Cognitive Load Theory also provides a plausible explanation for the empirical finding that users tend to expect stricter quality in longer documents, even when the overall error density remains stable [waddington1999, waddington2001]. This reflects a protective mechanism: anticipating that accumulated disruptions will erode comprehension, readers subconsciously raise the standard for what they consider acceptable in large-scale content. Together, these cognitive effects mirror and reinforce the non-linear sensitivity modeled by the Weber–Fechner framework. Although psychophysics focuses on the perceptual scaling of individual stimuli, Cognitive Load Theory models the impact of cumulative disruption. Both lead to the same conclusion: the relationship between perceived translation quality and the raw error count (that is, the objective number of penalty points) is fundamentally non-linear.
@@ 4.3 Perception and Cognition Converge on Logarithmic Scoring
In summary, classical psychophysical scaling [fechner1860, stevens1957] and cognitive load mechanics [sweller1988, miller1956] converge in the same functional form: tolerance for errors grows rapidly at first, then slows logarithmically. A calibrated logarithmic curve therefore offers the most parsimonious way to reconcile (i) diminishing marginal impact of individual errors with (ii) the accelerating cognitive cost of cumulative errors. This synthesis underpins the non-linear scoring model proposed in the next section.
@@ 5 Proposed Non-Linear Scoring Model
This paper adopts the terminology introduced in or primary study [GladkoffEtAl2024], distinguishing between three classes of scoring models.
Raw Score: $100-\alpha\,\mathrm{APT}/\mathrm{EWC}$, with $\alpha$ a unit factor (e.g., $\alpha=1000$ for “per 1,000 words”). This uses the sample’s penalty rate and is not a calibrated/anchored score. Calibrated Linear: This scoring model maps a user/institution’s acceptance range (pass/fail tolerance) onto a common 0–100 quality scale, making results easy to read and compare across jobs. Calibrated Non-Linear: A model that replaces the straight-line rule with a logarithmic tolerance curve. Relative to a linear rule anchored at the reference size, the model is more lenient for short samples and stricter for long ones; total allowance still grows with length, but at a slowing rate.
@@
Therefore, we recommend the non-linear, logarithmic model for all cases where evaluation sample sizes are expected to vary by more than $20\%$ from the reference wordcount (i.e., outside the $\pm 20\%$ fidelity interval defined in Section 5.4). Both the traditional raw model and the calibrated linear model remain viable tools, but their use should be restricted to scenarios where evaluation sample sizes consistently hover very close to the specific word count for which the models were calibrated. 666‘Calibration’ for the raw score simply means choosing the user’s pass/fail cut-point on the raw penalty-rate scale. Let $r_{\text{thr}}$ be the maximum allowed error rate (e.g., 5 points per 1,000 words). A sample with penalty total $\mathrm{APT}$ and word count $\mathrm{EWC}$ passes if and only if $\mathrm{APT}/\mathrm{EWC}\leq r_{\text{thr}}$, equivalently $\mathrm{APT}\leq r_{\text{thr}}\cdot\mathrm{EWC}$. No calibration mapping to a 0–100 quality scale is involved. Note that some teams report a 0–100 “raw score” by first scaling the penalty rate to a per-1,000-word unit and then computing $100-1000\cdot\mathrm{APT}/\mathrm{EWC}$; however, this rescaling still inherits the limitations of raw scoring. To obtain comparable scores across evaluation sample sizes, a calibrated model (linear near the reference size or non-linear for wider ranges) is required.
This section introduces the mathematical form of the proposed logarithmic model, explains how it can be calibrated, and demonstrates its practical advantages. By introducing a Non-Linear Calibrated scoring model, this paper extends the MQM scoring apparatus itself, enabling it to represent perceptual tolerance that linear models cannot capture. Because the same logarithmic function adapts seamlessly across evaluation sample sizes, language pairs, and content domains, an MQM-based Non-Linear Calibrated metric becomes congruent with holistic (document-level) expert judgment—turning MQM into a practically universal translation quality evaluation system.
5.1 Review the Limitations of Linear Extrapolation
@@
Linear scoring models typically assume that acceptable error counts increase proportionally with the sample size. For example, if a user accepts 5 minor errors in a 1000-word sample, the model allows 10 errors in 2000 words, 15 in 3000, etc. However, as shown in Sections 3 and 4, this assumption may be misaligned with the way content users perceive quality.
@@
As noted above, surveyed institutional users describe their scoring guidelines in terms that sound linear—roughly “about one serious error per page” as a tolerable upper bound. In practice, however, evaluators do not extend that allowance proportionally. When the sample expands to seven pages, they do not simply treat seven serious errors as acceptable; instead, reviewers report that they would typically fail such a sample already at three or four serious errors. In other words, tolerance tightens as the sample grows: the acceptable error density falls with text length. What matters is not a linear accumulation of defects, but how those defects shape the perceived quality of the translation as a whole. This asymmetry shows why linear extrapolation is misleading: it inflates the acceptable error count in longer samples and diminishes the perceived severity of errors in shorter ones.
5.2 Proposed Mathematical Model of Acceptable Errors
@@
Throughout, $\ln$ denotes the natural logarithm. If $x$ is measured in pages with $W$ words/page, replace $b$ by WbW\,b; $a$ is unchanged. This unit invariance ensures figures shown in pages and formulas written in words are consistent. The curve satisfies $E(0)=0$, rises steeply for short texts, and then flattens. This curve shape effectively combines both theoretical expectations and observed data.
To clarify the functional form, we adopt the intercept-free specification $E(x)=a\,\ln(1+bx)$. This choice enforces the natural boundary condition $E(0)=0$ and provides explicit control over curvature through the parameter $b$. This formulation is distinct from the “logarithmic trendline” $c+k\ln x$ commonly implemented in spreadsheet software, which introduces a free intercept term and does not pass through the origin. Consequently, the latter cannot be anchored to calibration points in an equivalent manner.  [^7] A related line of work is the LOGIPEM/HOPE metric we developed for comparing MT engines in cases where human reviewers perceive clear quality differences that standard automatic metrics fail to capture [gladkoff-han-2022-hope]. LOGIPEM/HOPE targets human-aligned system ranking at the document level by shaping how evidence from errors and discourse phenomena is aggregated; it is therefore complementary to the present contribution. In particular, LOGIPEM/HOPE addresses what is measured for engine comparison, whereas the non-linear tolerance function $E(x)$ proposed here addresses how pass/fail tolerance should vary with evaluation size. The former improves agreement with expert judgments across systems; the latter keeps scores consistent and comparable as sample length varies.
Other scoring models may be devised for specific applications. Importantly, however, an exponential ladder of severity weights is not the same thing as a non-linear scoring model in the sense used here. Severity weights (even exponential ones) change how individual errors are combined within a fixed sample; they do not by themselves make the pass/fail tolerance as a function of sample size non-linear.
For example, the ATA certification rubric is an analytic, points-off scheme tailored to short, fixed-length exam translations. Its severity scale is steep (many exponential levels), so higher-severity errors count far more than minor ones, but aggregation is still essentially additive and the exam length is fixed; there is no tolerance curve $E(x)$ that varies with $x$. Similarly, SAE J2450 (automotive) assigns fixed category weights, producing a non-uniform penalty distribution but still summing points linearly within a sample [saeJ2450]. (One could reproduce similar effects by choosing MQM type weights accordingly.) Thus ATA and SAE J2450 introduce severity-level non-uniformities, not the sample-size non-linearity addressed in this paper. They are effective for their narrow contexts, but they do not provide the general, non-linear tolerance function $E(x)$ needed to keep scores comparable when evaluation size varies widely.
@@ 5.3 Anchoring and Calibration Procedures
  • Anchoring defines the maximum acceptable error limit for a standard amount of text. This tolerance is expressed in penalty points based on a user’s real-world standard, such as, “no more than 5 penalty points per 1,000 words.” (one minor error is typically equivalent to one penalty point).
@@
Feasibility (two‑point).
Let $(x_{0},E_{0})$ and $(x_{1},E_{1})$ be two tolerance points with $x_{0}\neq x_{1}$ and $E_{0},E_{1}>0$. Define
@@
  • A second, empirical data point: This point is derived from observing the user’s real-world judgment (e.g., through a survey), such as their intuitive tolerance for a different sample size—for instance, accepting only 2 errors ($E_{1}=2$) in a 250-word text ($x_{1}=250$).
@@
See  Calibration from Two Points (Numerical) for a one-dimensional root-finding procedure in $b$ (e.g., bisection), followed by $a=E_{0}/\ln(1+b\,x_{0})$.
@@ 5.5 Visual comparison across three sample-size ranges
Figure 5 visualizes this critical weakness of the linear model: Once the slope is fixed using a single calibration point, the linear model severely diverges from the actual human judgment at both lower and higher evaluation sample sizes. In contrast, the logarithmic model aligns more closely with observed data regardless of sample size (evaluation word-count).
The calibrated logarithmic curve is more permissive for short samples and stricter for long ones relative to an anchored linear rule; it increases with length but flattens, avoiding the linear model’s over‑leniency at large sizes. This not only matches human content users’ expectations, but also reduces misclassification of borderline quality outputs.
@@
In summary, building on the Multi-Range Framework [GladkoffEtAl2024], this paper distinguishes three regimes of evaluation sample sizes based on their statistical and perceptual characteristics. In the micro-range (fewer than 250 words), evaluation outcomes are highly sensitive to single errors and statistical confidence is very low; thus, risk-based Statistical Quality Control (SQC) is most appropriate.
@@
Ultimately, however, for any operational scenario involving a wide variety of evaluated text lengths, the non-linear logarithmic model is the only option that ensures consistent and reliable accuracy. This is because a single linear extrapolation becomes unreliable as it moves further from its anchor point, whereas a logarithmic model remains accurate across both the meso- and macro-ranges. Figure 6 summarizes these applicability ranges, illustrating how tolerance curves change shape across evaluation sample size regimes.
@@ 6 Practical Implications
To the extent that error tolerance is not linear — as suggested by both empirical findings and practitioner reports — the adoption of a logarithmic scoring model can reshape translation quality evaluation practices, quality management processes, and software tooling across the industry. Replacing the linearity assumption allows quality assessment to become more accurate, consistent, and aligned with human perception. Looking ahead, such a model might also serve as a reference framework for emerging AI-based document-level evaluations. Current AI systems often provide holistic but qualitative judgments with uncertain completeness; a calibrated non-linear scoring model could, in principle, offer a quantitative reference point for future systems that integrate or adjust their evaluations against established calibrated 0–100 quality scale (user-specific PT/MSV mapping).
@@ 7 Building your non-linear scorecard
We list the following steps in this section including model and least-squares fit, from tolerance to a calibrated score, and scorecard mapping, followed by implications for CAT tools and linguistic quality assessment (LQA) automation, improvement of a unified calibrated quality scale, applications beyond human translation, and document-level evaluation. [^9] The designations and symbols in this section follow the variable naming used in the ASTM working item WK46396 (“MQM 2.0: Analytic Translation Quality Evaluation”), which describes the MQM scoring models and their mechanics in detail. http://www.astm.org/workitem-wk46396
@@ 7.3 Scorecard mapping
This keeps the MQM table (error types, error weights, severities and related multipliers $\rightarrow$ APT) unchanged; only the tolerance is dynamic via Equation  (2) and the calibrated mapping stays the same as the linear model’s PT/DPI (Passing Threshold/Defined Passing Interval) presentation.
Figure 7: Example of a real-world non-linear calibrated scoring model for documentation content type (solid yellow curve) plotted against the linear model with 10 allowed penalty points per 1000 words (dashed line). The logarithmic curve is calibrated from two anchor points (not shown in this plot), and therefore represents the actual user tolerance function; the linear line is drawn for comparison only.